OpenStudy (anonymous):

What is the value of the coefficient of "a" when the quadratic equation y = (2x - 1)(4x - 3) is written in standard form? Answer 6 8 -10 -6

OpenStudy (anonymous):

1.) y = (2x + 1)(4x - 3) y = 8x² - 6x + 4x - 3 y = 8x² - 2x - 3 y = 8(x² - 1/4 x - 3/8) The coefficient is 8.

OpenStudy (anonymous):

hope this helps Brad :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw